CH/DOC/94/10

Publications

Summary of Current Status

Author: B.R. Fulton

Date: 21 April 1994

This will continue to be a regular agenda item and I would
like to see the status of each experiment reviewed at each meeting.
Experiment   Anal  Comment                 Status                 Action
23Na data    NSJ   PRC paper               Ref report received    DLW
16O target   CDJ   Proposed PRC                                   CDJ
Texas data   JTM   PRC/NPA                 1st draft circulating  BRF
Higher Ex    MJL   PRC/NPA                 1st draft circulating  MJL
LBL 16O+12C  JTM   PRC/NPA                 1st draft circulating  BRF
Leddy Det    MJL   NIM's article           Started 1st draft      MJL
Hybrid       NC/SMS Nim's article          1st draft near         NMC/WNC
Band term.   SPF   Brief report (see next)                        DLW
Spin low E   CDJ   Overlaps GJG paper; merge with previous?       CDJ
24Mg+24Mg    AStJM Aim to set limit on 16O+16O+16O                on hold
7 alpha      PMS   Analysis still uncertainties                   WDMR
12C on 24Mg  SMS   Sort O+Be channel; new data from ANU to sort   WNC
LBL 16O+8Be  JSP   Analysis nearly at stage to draft paper        BRF
High Res.    NC    Analysis done. Monte Carlo done. Start draft   NMC
12C+12C*     RPW   Analysis needs 12 months                       RPW
Ang. Dist.   SPC   Recoil ambiguities + poor stats. Analysis cont DLW
40Ca         SMS   In analysis                                    WNC
28Si/32S     NC    In analysis                                    NMC
7-alpha      PS    (see above)                                    WDMR
7-alpha      MF    Probable paper on Be-Be states                 MF
03+          SPC   In analysis                                    DLW
C+C Ex Fn    CDJ   In analysis                                    CDJ
C+O Ex Fn    SPF   In analysis                                    DLW
O+C          GJG   In analysis                                    GJG
Megha Det    PRL   In analysis  Hope for NIM's paper              DLW